Two friends have built a lemonade stand, and they plan to start out by selling 20 liters of lemonade containing 23% lemon juice. To make this, they will mix some homemade lemonade containing 27% lemon juice with some store-bought lemonade that contains 22% lemon juice. How much of each type should the friends use?
Answer by amalm06(224)

The problem can be solved in two different ways. Consider first the following system of equations:

x+y=20

0.27x+0.22y=(0.23)(20)

Solving the above equations for x and y will give you the relevant quantities.

The other method is that of alligation.

Let S1 denote the 22% solution and S2 denote the 27% solution.

Then 23-22=1 and 27-23=4. So that S1:S2=4/1. This mean means that we have 4 parts of S1 for every 5 parts of mixture. We can find S1 and S2 as follows:

S1=(4/5)(20) = 16 L (Answer)

S2=20-16=4 L (Answer)